A 4.8 cm × 5.8 cm rectangle lies in the xy-plane.

a) What is the electric flux through the rectangle if E⃗ =(110ı^−220k^)N/C?

b) What is the electric flux through the rectangle if E⃗ =(110ı^−220ȷ^)N/C?
Answer #1

Anea = 4.80 5.8x10-4 m² = 27984x10 vector is toris Ê » a) 0 - E-cis = (lloi -220 €) · (27-84x104 €) =-0 - 220x27-84 8 10 4k »